OpenStudy (anonymous):

How would you solve 1 + 10-2/4 But the 10-2 is above the 4 like a fraction?

hartnn (hartnn):

like this ? \(1+\dfrac{10-2}{4}\) ? first calculate whats 10-4 =... ?

hartnn (hartnn):

****10-2 =... ?

OpenStudy (anonymous):

Yesh

OpenStudy (anonymous):

How do you do that ?!

hartnn (hartnn):

how did i do what ? write like that ?

OpenStudy (anonymous):

Write the problem like that?

hartnn (hartnn):

we can either use \(Latex\) or equation editor (\(\sum Equation\)) provided below using latex, the code is `\( 1+\dfrac{10-2}{4} \)`

hartnn (hartnn):

which gave me \(1+\dfrac{10-2}{4}\)

OpenStudy (texaschic101):

you having trouble with this problem ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

first calculate whats 10-2 = .... ?